Ariana took 2h longer to drive 360mi on the first day of a trip than she took to drive 270mi on the second day. If her speed was the same on both days, what was the driving time each day?
Thank you for helping me understand these problems:-)

Answer by checkley71(8403)

DISTANCE=RATE*TIME
360=R(T+2)
270=RT SOLVING FOR R IN THIS EQUATION WE GET
R=270/T NOW SUBSTITUTING (270/T) FOR R IN THE FIRST EQUATION WE GET
360=(270/T)(T+2)
360=(270T+540)/T
360T=270T+540
360T-270T=540
90T=540
T=540/90
T=6 HOURS IS THE TIME FOR THE RETURN 270 MILE TRIP.
6+2=8 HOURS FOR THE FIRST TRIP WHICH TOOK 2 HOURS LONGER.
THE RATE IS
270/6=45 MPH
###########################################################
MERRY CHRISTMAS